OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = 3x + 4. What is f(g(x)) ?

OpenStudy (anonymous):

I believe it should be like this. x-2(3x + 4) now use the foil method = 3x^2 + 4x - 6x -8 = 3x^2 -2x - 8

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

and to solve further if u want to find what the value of f, u will factorize 3x^2 - 2x -8 (3x + 4) (x-2) 3x + 4 = 0 or x-2=0 3x = -4 x = -4/3 or x=2